# Smooth Tchebycheff Scalarization for Multi-Objective Optimization

### v2

@@ Conclusion
This work only analyzes the properties of STCH scalarization for deterministic multi-objective optimization. It could be interesting to investigate STCH scalarization for stochastic optimization in future work.
@@ Appendix A Detailed Proofs
$\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\hat{}\bm{x})-z^{*}_{i})}/{\mu}}\right){\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}\leq}\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\bm{x}^{*})-z^{*}_{i})}/{\mu}}\right).$ (16)
@@
[…]
@@
Appendix B Practical Implementation
@@ B.3 Setting the Smoothing Parameter $\mu$
Appendix C Detailed Experiment Setting
C.1 Multi-Task Learning Problems
@@
In this work, we implement our STCH scalarization method for MTL problems with the LibMTL library (Lin & Zhang, 2023), and follow the experiment settings of Lin et al. (2023).
@@
NYUv2 (Silberman et al., 2012) is an indoor scene understanding dataset with $3$ tasks on semantic segmentation, depth estimation, and surface normal prediction, with $795$ training and $654$ testing images. Following Lin et al. (2023), we compare all MTL methods using the SegNet model with a shared encoder and 3 task-specific decoders. The model is trained for $200$ epochs with Adam (Kingma & Ba, 2015), of which the learning rate is initially set to $10^{-4}$ with $10^{-5}$ weight decay and will be halved to $55\times 10^{-}$ 5 after $100$ epochs. The batch size is set to $2$. The three optimization objectives are cross-entropy loss, $L_{1}$ loss and consine loss for semantic segmentation, depth estimation, and surface normal prediction.
@@
QM9 (Ramakrishnan et al., 2014) is a molecular property prediction dataset with 11 tasks on different properties. The data split in Navon et al. (2022) is used to devide the dataset into $110,000$ for training, $10,000$ for validation, and $10,000$ for testing. All MTL methods are compared on the same model with a shared graph neural network encoder and $11$ linear task-specific head as in Navon et al. (2022) and Lin et al. (2023). All methods are trained by Adam (Kingma & Ba, 2015), of which the leanring rate is $10^{-3}$ with the ReduceLROnPlateau scheduler. The batch size is $128$ and the number of training epochs is $300$. The optimization objectives are $11$ mean squared error (MSE) for each task.
C.2 Pareto Set Learning
@@
C.2.1 Synthetic Benchmark Problems
@@
F1
$f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\sqrt{\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}}\right)$
where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-(2\bm{x}_{1}-1)^{2})^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-(2\bm{x}_{1}-1)^{2})^{2},$
$J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (43)
F2
$f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\sqrt{\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}}\right)$
where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-\bm{x}_{1}^{0.5(1.0+\frac{3(j-2)}{n-2})})^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-\bm{x}_{1}^{0.5(1.0+\frac{3(j-2)}{n-2})})^{2},$
$J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (44)
F3
$f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\sqrt{\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}}\right)$
where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-\sin(4\pi\bm{x}_{1}+\frac{j\pi}{n}))^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-\sin(4\pi\bm{x}_{1}+\frac{j\pi}{n}))^{2},$
$J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (45)
F4
$f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\left(\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}\right)^{2}\right)$
where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-(2\bm{x}_{1}-1)^{2})^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-(2\bm{x}_{1}-1)^{2})^{2},$
$J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (46)
F5
$f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\left(\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}\right)^{2}\right)$
where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-\bm{x}_{1}^{0.5(1.0+\frac{3(j-2)}{n-2})})^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-\bm{x}_{1}^{0.5(1.0+\frac{3(j-2)}{n-2})})^{2},$
$J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (47)
F6
$f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\left(\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}\right)^{2}\right)$
where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-\sin(4\pi\bm{x}_{1}+\frac{j\pi}{n}))^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-\sin(4\pi\bm{x}_{1}+\frac{j\pi}{n}))^{2},$
$J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (48)
C.2.2 Real-World Engineering Design Problems

### v3

@@ Conclusion
This work only focuses on unconstrained and deterministic multi-objective optimization. It could be interesting to investigate how to properly handle different constraints and tackle stochastic optimization with STCH scalarization in future work. A brief discussion on these directions can be found in Appendix B.4 and B.5.
Acknowledgements
The work described in this paper was supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (GRF Project No. CityU11215622), the National Natural Science Foundation of China (Grant No. 62106096), the Natural Science Foundation of Guangdong Province (Grant No. 2024A1515011759), the National Natural Science Foundation of Shenzhen (Grant No. JCYJ20220530113013031).
@@ Appendix A Detailed Proofs
$\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\hat{}\bm{x})-z^{*}_{i})}/{\mu}}\right){\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\leq}\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\bm{x}^{*})-z^{*}_{i})}/{\mu}}\right).$ (16)
@@
Theorem 3.10 (Convergence to Pareto Stationary Solution). If there exists a solution ^$\hat{}\bm{x}$ such that $\nabla g^{(\text{STCH})}_{\mu}(\hat{}\bm{x}|\bm{\lambda})=0$, then ^$\hat{}\bm{x}$ is a Pareto stationary solution of the original multi-objective optimization problem (1).
@@
Appendix B Practical Implementation for STCH Scalarization
@@ B.3 Setting the Smoothing Parameter $\mu$
B.4 Constrains Handling
This work mainly focuses on the unconstrained multiobjective optimization problem, while the constrained problem could also be important for many real-world problems. Extra approaches could be needed for STCH to handle the constraints.
A constrained multiobjective optimization problem can be defined as
$\min_{x\in\Omega}\ f(x)=(f_{1}(x),f_{2}(x),\cdots,f_{m}(x))$ (42)
with $\Omega=\{x\in R^{n}:g_{j}(x)\leq 0,j=1,\ldots,p,\quad h_{l}(x)=0,l=1,\dots,q\}$, where $\{g_{j}(x)\}$ and $\{h_{l}(x)\}$ are the inequality and equality constraints. For differentiable multiobjective optimization, we assume that all objectives $\{f_{i}(x)\}$ and constraints $\{g_{j}(x)\},\{h_{l}(x)\}$ are continuously differentiable.
Following previous work on gradient-based multiobjective optimization (Fliege & Vaz, 2016), we can reformulate the constrained multiobjective optimization problem into an unconstrained problem with additional objectives:
$\min_{x\in R^{n}}\ (f(x),g^{+}(x),|h(x)|)=(f_{1}(x),\cdots,f_{m}(x),g^{+}_{1}(x),\cdots,g^{+}_{p}(x),|h_{1}(x)|,\cdots,|h_{q}(x)|)$ (43)
where $g^{+}_{j}(x)=\max\{0,g_{j}(x)\}$. A feasible solution $x$ for the original constrained problem satisfies $\{g^{+}_{j}(x)=0,j=1,\ldots,p,\quad|h_{l}(x)|=0,l=1,\dots,q\}$. Instead of the original problem with $m$ objectives and $p+q$ constraints, we now have an unconstrained problem with $m+p+q$ objectives and 00 constraints. With our proposed STCH scalarization, we have the following problem:
$g^{(\text{STCH})}_{\mu}(x|\lambda,\sigma)=\mu\log\left(\sum_{i=1}^{m}e^{\frac{\lambda_{i}(f_{i}(x)-z^{*}_{i})}{\mu}}+\sum_{j=1}^{p}e^{\frac{\sigma g^{+}_{j}(x)}{\mu}}+\sum_{l=1}^{q}e^{\frac{\sigma|h_{l}(x)|}{\mu}}\right),$ (44)
where $\sigma\geq 0$ is a penalty parameter. We can also use the smooth counterparts for $g^{+}_{j}(x)$ and $|h_{l}(x)|$ in the STCH scalarization above to achieve a better convergence rate. In addition to this soft constraint handling approach, there could be many other possible constraint handling methods that work for STCH, which is an important research direction in the future.
B.5 Potential Extension for Stochastic Multi-Objective Optimization
In this work, we mainly focus on deterministic multi-objective optimization setting, and leave the study for stochastic optimization as an important future work. We would like to make the following remarks on this concern:
  • Stochastic multi-objective optimization itself is an important research topic that was not yet well explored in the past decades. Most existing methods only study the deterministic setting. Until recently, a few work have been proposed to investigate the stochastic adaptive gradient method (Liu & Vicente, 2021; Zhou et al., 2022; Fernando et al., 2023). In this work, instead of yet another adaptive gradient method, we focus on proposing a lightweight, efficient, and theoretically solid scalarization method.
  • Although developed from the deterministic setting, our proposed STCH scalarization method can outperform MOCO (Fernando et al., 2023), a state-of-the-art stochastic adaptive gradient method on all multi-task learning datasets (for performance and run time), which are essentially stochastic optimization problems. These results validate that our proposed method can be used to tackle stochastic optimization problems.
  • In addition, how to design an efficient algorithm for stochastic nonsmooth optimization problem is also an emerging research topic. A recent work (Wang & Zhang, 2023) investigates efficient accelerated gradient methods for the classic smoothing method proposed in Nesterov (2005) and Beck & Teboulle (2012). We might leverage this recent development to investigate efficient scalarization methods for stochastic multi-objective optimization.
Appendix C Discussion with Related Work on Minimax Optimization
The Tchebycheff scalarization approach is naturally a minimax optimization problem. In this section, we briefly discuss the relation between our proposed STCH method and the related algorithms for minimax optimization.
Problem Formulation
A general minimax optimization problem is in the form:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}f(x,y),$
where $\mathcal{X}\in R^{n}$ and $\mathcal{Y}\in R^{m}$ are both closed and bounded convex sets, while $f:\mathcal{X}\times\mathcal{Y}\rightarrow R$ is a smooth function. From the viewpoint of Tchebycheff scalarization, we are more interested in the problem with a specific linear structure on $y$:
$\min_{x\in\mathcal{X}}\max_{y\in\Delta^{m-1}}F(x)^{T}y,$
where $F(x)=(f_{1}(x),f_{2}(x),\cdots,f_{m}(x))^{T}$ and $\Delta^{m-1}=\{y|\sum_{i=1}^{m}y_{i}=1,y_{i}\geq 0\ \forall i\}$ is a probability simplex. It is easy to check that the optimal solution of the above problem is also optimal for the Tchebycheff scalarization problem $\min_{x\in\mathcal{X}}\max_{1\leq i\leq m}f_{i}(x)$ since the linear optimization problem over the simplex is always optimized at one of its vertices (Boyd & Vandenberghe, 2004). Here we drop the preference $\lambda$ and reference point $z^{*}$ for better exposition.
Algorithms for Minimax Optimization
There are many algorithms to solve the minimax optimization problems, such as some classic methods (Nemirovski, 2004; Nesterov, 2007; Monteiro & Svaiter, 2010). In recent years, different algorithms have been proposed to tackle the minimax problems in different settings (Palaniappan & Bach, 2016; Gidel et al., 2017; Mertikopoulos et al., 2019). Rather than those algorithms with a double-loop or even triple-loop structure, we are more interested in the lightweight single-loop algorithm for minimax optimization such as Gradient Descent Ascent (GDA) (Lin et al., 2020a). However, the simple GDA algorithm could oscillate during the optimization process (Mokhtari et al., 2020) and fail to converge even for a simple bilinear problem $\min_{x\in R^{n}}\max_{y\in R^{n}}x^{T}y$ (Letcher et al., 2019). Extra efforts, such as alternating gradient projection (Xu et al., 2023) or a smoothing method (Zhang et al., 2020), are needed for GDA.
Comparison and Discussion
The smoothed gradient descent ascent (smoothed-GDA) (Zhang et al., 2020) algorithm is the most related minmax optimization algorithm to our proposed STCH method. The key approach in smoothed-GDA is to add an extra smoothed regularization term to the objective function:
$K(x,z;y)=f(x,y)+\frac{p}{2}||x-z||^{2}.$ (45)
To optimize the above function, in addition to the original $x$ and $y$, smoothed-GDA maintains and updates an auxiliary sequence $\{z^{t}\}$ where $z^{t+1}=z^{t}+\beta(x^{t+1}-z^{t})$ throughout the optimization process. The terms $p>0$ and $\beta>0$ are hyperparameters, and the smoothed-GDA reduces to the standard GDA when $\beta=1$.
Compared to smoothed-GDA, our proposed STCH approach does not require maintaining an extra sequence $\{z^{t}\}$ which could lead to high memory cost (e.g., an additional deep neural network with millions of parameters). For Pareto set learning $x^{*}(\lambda)=h_{\theta}(\lambda)=argmin_{x\in\mathcal{X}}g(x|\lambda)\forall\lambda\in\Delta^{{m-1}}$, we need to randomly sample different trade-off preferences $\lambda\in\Delta^{{m-1}}$ (and hence different single-objective subproblems) at each iteration. In this case, it could be very hard to maintain extra sequences for each subproblem, since the trade-offs could be infinite and come in an online manner. Our proposed STCH method is simple, lightweight, and with good theoretical properties, and we hope that it can inspire more follow-up work on developing more efficient gradient-based methods (e.g., improved minmax optimization algorithms) for multi-objective optimization.
$\mu$ $\alpha$ $\epsilon_{HOMO}$ $\epsilon_{LUMO}$ $<R^{2}>$ ZPVE $U_{0}$ $U$ $H$ $G$ $c_{v}$ $\Delta_{p}\uparrow$
TCH 0.266 0.401 107.1 151.6 5.922 13.2 166.7 167.5 168.1 162.0 0.206 -252.2
Smoothed-GDA 0.252 0.424 105.2 157.1 4.561 11.8 162.4 159.3 149.9 152.2 0.198 -218.8
STCH (Ours) 0.166 0.260 94.48 101.2 1.850 4.88 58.34 58.68 58.70 58.27 0.104 -58.14
We also implemented and tested smoothed-GDA on the QM9 dataset. According to the results above, smoothed-GDA achieves better overall performance than the original TCH scalarization, but is significantly outperformed by our proposed STCH scalarization method. We have also tried to tune the hyperparameters for smoothed-GDA, but do not obtain any significant improvement.
Appendix D Detailed Experiment Setting
D.1 Multi-Task Learning Problems
@@
In this work, we implement our STCH scalarization method for MTL problems with the LibMTL library (Lin & Zhang, 2023), and follow the experiment settings of Lin et al. (2023a).
@@
NYUv2 (Silberman et al., 2012) is an indoor scene understanding dataset with $3$ tasks on semantic segmentation, depth estimation, and surface normal prediction, with $795$ training and $654$ testing images. Following Lin et al. (2023a), we compare all MTL methods using the SegNet model with a shared encoder and 3 task-specific decoders. The model is trained for $200$ epochs with Adam (Kingma & Ba, 2015), of which the learning rate is initially set to $10^{-4}$ with $10^{-5}$ weight decay and will be halved to $5\times 10^{-5}$ after $100$ epochs. The batch size is set to $2$. The three optimization objectives are cross-entropy loss, $L_{1}$ loss and consine loss for semantic segmentation, depth estimation, and surface normal prediction.
@@
QM9 (Ramakrishnan et al., 2014) is a molecular property prediction dataset with 11 tasks on different properties. The data split in Navon et al. (2022) is used to divide the dataset into $110,000$ for training, $10,000$ for validation, and $10,000$ for testing. All MTL methods are compared on the same model with a shared graph neural network encoder and $11$ linear task-specific head as in Navon et al. (2022) and Lin et al. (2023a). All methods are trained by Adam (Kingma & Ba, 2015), of which the learning rate is $10^{-3}$ with the ReduceLROnPlateau scheduler. The batch size is $128$ and the number of training epochs is $300$. The optimization objectives are $11$ mean squared error (MSE) for each task.
D.2 Pareto Set Learning
@@
D.2.1 Synthetic Benchmark Problems
@@
F1 $f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\sqrt{\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}}\right)$ where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-(2\bm{x}_{1}-1)^{2})^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-(2\bm{x}_{1}-1)^{2})^{2},$ $J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (47)
F2 $f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\sqrt{\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}}\right)$ where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-\bm{x}_{1}^{0.5(1.0+\frac{3(j-2)}{n-2})})^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-\bm{x}_{1}^{0.5(1.0+\frac{3(j-2)}{n-2})})^{2},$ $J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (48)
F3 $f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\sqrt{\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}}\right)$ where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-\sin(4\pi\bm{x}_{1}+\frac{j\pi}{n}))^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-\sin(4\pi\bm{x}_{1}+\frac{j\pi}{n}))^{2},$ $J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (49)
F4 $f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\left(\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}\right)^{2}\right)$ where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-(2\bm{x}_{1}-1)^{2})^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-(2\bm{x}_{1}-1)^{2})^{2},$ $J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (50)
F5 $f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\left(\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}\right)^{2}\right)$ where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-\bm{x}_{1}^{0.5(1.0+\frac{3(j-2)}{n-2})})^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-\bm{x}_{1}^{0.5(1.0+\frac{3(j-2)}{n-2})})^{2},$ $J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (51)
F6 $f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\left(\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}\right)^{2}\right)$ where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-\sin(4\pi\bm{x}_{1}+\frac{j\pi}{n}))^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-\sin(4\pi\bm{x}_{1}+\frac{j\pi}{n}))^{2},$ $J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (52)
D.2.2 Real-World Engineering Design Problems